ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: Electronic Publishing Buswell, Stephen, Stilo Technology Ltd An Introduction to MathML Mathematical Markup Language (MathML) has recently been published as a Recommendation by the World Wide Consortium (W3C). The goal of MathML is to enable mathematics to be served, received and processed on the Web. This paper discusses the issues involved with display and processing of mathematics on the Web and outlines the XML-based solution developed. This solution supports the encoding of both presentation and semantic content of mathematical expressions. The relationship between MathML and existing standards and formats such as TEX is discussed. An overview of MathML-compliant tools (existing and under development) for the creation and processing will be given.